OpenStudy (anonymous):

3(-8)-l4(-2)l simplify

OpenStudy (anonymous):

i got -24+294=270

OpenStudy (anonymous):

3(-8)-|4(-2)| -24-|-8| -24-8 -32

OpenStudy (anonymous):

Is that a 1 in front of the 4 or an absolute value bar?

OpenStudy (anonymous):

Not quite. \(3(-8) = -24\) That part was right. So \(3(-8) - |4(-2)|\) \[=-24 - |-8|\]\[=-24 - (8)\]\[= -24-8\]\[= -32\]